The relation between time t and distance x is t = ax2 + bx where a and b are constants. The acceleration is ?

Grade:12

1 Answers

dev bhatnagar
15 years ago

t = ax2 + bx

=> Differentiate both side with respect to time (t)

=> dt/dt = 2ax.dx/dt + b.dx/dt

=> 1 = 2axv + bv

=> v.(2ax + b) = 1

=> (2ax + b) = 1/v

=> Again differentiate both side with respect to time (t)

=> 2a.dx/dt = -v-2 . dv/dt

=> 2av = v-2 . acceleration

=> acceleration = -2av3

Thanks